# Curriculum Approximate Unlearning for Session-based Recommendation

### v1

@@ 4.1. Multi-objective unlearning
While effective unlearning is important, maintaining recommendation performance is equally crucial. This necessitates a careful balance between the unlearning objective and the recommendation utility, so we formulate the approximate unlearning as a multi-task optimization problem. The key idea is to jointly optimize the recommendation utility and unlearning effectiveness through training a composite objective function with three task-specific losses: (1) the unlearning loss $\mathcal{L}_{unlearn}$ that encourages forgetting specific samples, (2) the retention loss $\mathcal{L}_{normal}$ to preserve recommendation performance, and (3) the consistency loss $\mathcal{L}_{KL}$ that regularizes deviations from the previous model.
Next, we will introduce the three types of loss functions in detail and describe how they are weighted during optimization.
Unlearn Loss. Unlearning can be regarded as the inverse process of forward learning, so we apply gradient ascent to explicitly remove the influence of the unlearning sample, which intuitively corresponds to adding back the gradient of the target data that was previously subtracted. For the target item $v_{t}$ to be unlearned, the unlearning loss is formulated as:
(5) $\mathcal{L}_{unlearn}=log(P_{\theta}(v_{t}|v_{1:t-1}))$
@@
Normal Loss. To ensure that the unlearned model retains its general predictive capability, we include a standard cross-entropy loss based on the next item $v_{t+1}$, given the prefix session $\{v_{1},...,v_{t-1}\}$:
(6) $\mathcal{L}_{normal}=-log(P_{\theta}(v_{t+1}|v_{1:t-1}))$
Since unlearning is performed to remove an item from a session, it is essential to ensure that the model maintains continuity on remaining items in the session. $\mathcal{L}_{normal}$ guides the model to remain functional in its primary task, i.e., predicting plausible future items based on user behavior, despite forgetting specific items.
@@
(7) $\mathcal{L}_{KL}=KL(P_{\theta^{rec}}(\cdot|v_{1:t-1})||P_{\theta}(\cdot|v_{1:t-1}))$
@@
(8) $\mathcal{L}_{app}=\alpha_{1}\mathcal{L}_{unlearn}+\alpha_{2}\mathcal{L}_{normal}+\alpha_{3}\mathcal{L}_{KL}$
@@ 4.2. Pareto-Optimality solver
A crucial challenge in multi-task optimization is balancing the importance of each objective. Manually tuning the weights $\alpha_{1}$, $\alpha_{2}$, and $\alpha_{3}$ is suboptimal and sensitive to data distribution and model scale. Furthermore, maintaining fixed hyperparameters throughout the entire training process can hinder the model’s ability to adapt to changes and negatively impact performance. To effectively solve this problem, we employ a Pareto-Otimality solver, which can reliably find a stable solution that balances the trade-offs between conflicting objectives. This approach ensures that the model simultaneously achieves effective unlearning while maintaining overall recommendation quality. Given a model parameterized by $\theta$ and a series of learning objective functions $\mathcal{L}_{unlearn},\mathcal{L}_{normal},\mathcal{L}_{KL}$, Pareto-Optimality is a state in which it is impossible to enhance one objective without compromising another. Formally, the Pareto-Optimality can be defined as:
@@
Formally, MGDA uses the Karush-Kuhn-Tucker (KKT) conditions (Bertsekas, 1997) to describe constraints of finding $\alpha_{1},\alpha_{2},\alpha_{3}$:
(10) $\min_{\alpha_{1},\alpha_{2},\alpha_{3}}\left\|\sum_{i=1}^{3}\alpha_{i}\mathbf{g}_{i}\right\|^{2}_{2}\quad\text{subject to}\quad\sum_{i=1}^{3}\alpha_{i}=1,\ \alpha_{i}\geq 0$
@@
After computing $\alpha_{i}$ for all tasks, we utilize the loss weighted by $\alpha_{i}$ to update the model parameters, thus identifying an optimization direction that preserves the performance of recommendation while improving the effectiveness of unlearning.
@@ 4.3. Curriculum unlearning
To further enhance effectiveness of approximate unlearning when handling multiple requests simultaneously, we propose a curriculum-based unlearning strategy that prioritizes forgetting easier samples first, progressively moving toward harder ones. The underlying assumption is that removing low-impact easier samples earlier helps stabilize training and facilitates optimization. The curriculum unlearning framework comprises two key components: (1) difficulty estimation and (2) sample scheduling.
@@ 4.3.1. Difficulty estimation
(11) $Dif_{g}(x)=-\cos\left(\nabla_{\theta}\mathcal{L}_{unlearn}(x),\ \nabla_{\theta}\mathcal{L}_{normal}(x)+\nabla_{\theta}\mathcal{L}_{KL}(x)\right)$
@@
Embedding unlearning difficulty. This metric is based on the intuition that samples which the model strongly favors are harder to be forgotten. Specifically, we compute the dot product between the session state representation and the embedding of the item to be forgotten. Given a session state representation $\mathbf{e}_{session}$ and the embedding of the target item $x$ to be forgotten $\mathbf{e}_{x}$, CAU computes:
(12) $Dif_{e}(x)=\langle\mathbf{e}_{session},\ \mathbf{e}_{x}\rangle$
Here, $\mathbf{e}_{session}$ is the encoded state representation of the session prefix (e.g., from an RNN or self-attention network), and $\mathbf{e}_{x}$ is the embedding of the target item $x$ to be forgotten. $\langle\rangle$ denotes dot product. A higher $Dif_{e}$ implies the model is more confident in predicting the unlearning target $x$ , making it harder to be removed. Hence, we treat samples with lower $Dif_{e}$ as easier ones and schedule them earlier in the unlearning process.
@@
4   calculate $Dif_{g}$ or $Dif_{e}$ according to Eq.(11) or Eq.(12);
5 $D_{f\text{-sorted}}\leftarrow\text{sort }D_{f}\text{ ascendingly }\text{by $Dif_{g}$ or $Dif_{e}$}$;
@@ 4.3.2. Sample scheduling
For example, when employing the gradient unlearning difficulty metric, the initial phase of each epoch concentrates on samples with high gradient alignment (i.e., lower $Dif_{g}$), indicating that their gradients align well and thus are easier to be optimized. As training progresses, more challenging samples with higher $Dif_{g}$ and more conflicting gradients are incrementally introduced into the training batches. Such curriculum progression reduces abrupt gradient conflicts and helps maintain a stable optimization trajectory.
@@
(13) $p_{x}(t)=\frac{(\exp((2t-1)\cdot(Dif(x)-\bar{Dif})))^{\tau}}{\sum_{j}(\exp((2t-1)\cdot(Dif(j)-\bar{Dif})))^{\tau}}$
@@ 5 $t\leftarrow step/T$;
7      calculate DifDif according to Eq.(11) or Eq.(12);
@@ 16 $\theta^{app}\leftarrow$ update model parameter according to $\mathcal{L}_{app}$;
where $Dif$ denotes the difficulty scores (i.e., either $Dif_{g}$ or $Dif_{e}$), and $\bar{Dif}$ is the mean difficulty score of all unlearning samples. The time variable $t$ is normalized to the interval $[0,1]$ based on the current training progress, calculated as t=current_step/total_stept=current\_step/total\_step. $\tau$ is the temperature coefficient.
@@ 5.1.1. Datasets and pre-processing
We conduct experiments on three public accessible datasets: MovieLen-1M [^1] https://grouplens.org/datasets/movielens/1m/, Amazon Beauty and Games [^2] https://jmcauley.ucsd.edu/data/amazon/. These datasets vary in scale, domain, and sparsity.
@@ 5.1.2. Evaluation protocols
Given that recommendation performance and unlearning effectiveness follow distinct numerical patterns, such as higher $NDCG@k$ and $Recall@k$ indicate better recommendation performance, while lower $Hit_{u}@k$ signifies better unlearning effectiveness. We introduce a new metric $U_{\beta}$ to facilitate a comprehensive evaluation of both aspects. Specifically, $U_{\beta}$ integrates $Recall@k$ and $Hit_{u}@k$ as:
@@ 5.1.3. Baselines
  • GRU4Rec (Hidasi et al., 2015) utilizes gated recurrent units (GRU) to model user interaction sequences which is an RNN-based sequential recommendation model.
@@
  • Retrain: retrain the whole model from scratch on the remaining dataset. It’s computationally expensive.
@@ 5.1.4. Implementation details
All methods are learned with the Adam optimizer (Kingma and Ba, 2014). The learning rate is set as 1e31e-3. The model input consists of the most recent 10 interacted items for Beauty, 20 for Games, and 50 for MovieLens-1M, with shorter sequences padded using a special padding token. The training batch size is set to 256, while the unlearning batch size is set to 128. Temperature coefficient $\tau$ is 2. The embedding size is fixed at 64. For the unlearning process in hard-sampling, the unlearning epoch is set to 200 for SASRec and BERT4Rec, and 100 for GRU4Rec. And for soft-sampling, the number of unlearning step is adjusted correspondingly. Each experiment is repeated 5 times, and the average performance is reported.
Table 2. Unlearning effectiveness and overall performance comparison of different methods on the three datasets. The U-score is computed by Recall@10 and $\text{Hit}_{u}@1$, with $\beta=10$ for Beauty and $\beta=3$ for Games and ML-1M. Boldface denotes the best score.
@@ 5.2. Recommendation performance
We list the recommendation performance comparison results in Table 1. We can observe that due to the removal of 10% of the training data, the recommendation performance declines across all baselines, which is an expected outcome. However, in the context of the unlearning task, our primary concern is the ability to preserve recommendation performance on the remaining samples rather than the unlearning effectiveness alone.
From the table, We observe that the recommendation performance of SISA and SRU drops markedly, possibly due to the partitioning of the dataset into disjoint shards, which substantially reduces correlations between sessions. Specifically, even with the incorporation of attention-based aggregation mechanisms in SRU, its effectiveness in mitigating performance degradation remains limited. Compared with Retrain, their performance drops by 63% and 32%, respectively. Such a drastic reduction is undesirable, as overly low recommendation performance would directly impair the user experience.
In contrast, CAU demonstrates significantly superior performance preservation, achieving results close to the Original model with a stable performance drop controlled within 8%. This advantage can be attributed to the introduction of the normal loss and KL loss during the unlearning model update, which suppresses excessive disturbance to the original representation distribution, thereby enabling effective unlearning of target samples while retaining the model’s recommendation capability as much as possible.
@@ 5.3. Unlearning effectiveness and efficiency
What’s more, we can observe that in most cases, soft-sampling is more effective than hard-sampling. For instance, on the ML-1M and BERT4Rec datasets using gradient unlearning difficulty, the U-score of soft-sampling is 0.6913 compared to 0.6711 for hard-sampling. Similarly, with embedding unlearning difficulty, soft-sampling consistently achieves higher scores than hard-sampling. This improvement is likely due to the increased robustness that soft-sampling provides throughout the unlearning procedure.
Table 3. Comparison of unlearning efficiency (minute [m]). Boldface denotes the best score.
@@ 5.4. Ablation study
In this part, we conduct ablation study to analyze the functionality of the main components and strategies of CAU (i.e., Pareto-Optimality solver (PS), curriculum unlearning (CU), and recommendation loss (RL), where RL consists of both the normal loss and the KL loss). Table 4 shows the performance of CAU and its variants on all three datasets using SASRec as the backbone model. We introduce the variants and analyze the effects respectively.
(1) Effect of Pareto-Optimality solver: We removed the Pareto-Optimality solver component and assigned an equal weight to each task, which means setting all $\alpha$ values to 0.33. The experimental results demonstrate the effectiveness of the Pareto-Optimality solver: by formulating the entire unlearning process as a multi-task optimization problem, the Pareto-Optimality solver iteratively searches for gradient update directions that do not degrade any individual sub-task, thereby achieving a better balance between recommendation performance and unlearning effectiveness.
@@
(3) Effect of recommendation loss: We remove both the normal loss and the KL loss, which means directly apply GA to unlearn data. From the results we can find that this will lead to catastrophic effects, as the model’s recommendation accuracy would almost completely collapse within fewer than five unlearning epochs, and the $Hit_{u}$ value would become meaningless. This observation highlights the critical role of the recommendation loss in preserving the model’s recommendation capability.
@@ 5.5.1. Impact of unlearning batch size
In this part, we investigate the impact of batch size on the unlearning performance during the forgetting process. Because the Pareto-Optimality solver estimates task conflicts and updates the model on a batch basis, the batch size plays a critical role in determining its effectiveness.
Here we choose batch size $\in\{32,64,128,256,512\}$ to conduct experiments. The figure 2 illustrates the performance of the three models on the Games dataset. We clearly observe that as the batch size increases, the U-score tends to decrease. This is because the Pareto-Optimality solver dynamically searches for the optimal gradient update on a per-batch basis. And a smaller batch size reduces conflicts among data samples within each batch, increasing the likelihood of identifying accurate update directions, thereby yielding higher U-scores, but may lead to longer training time. We also measure the training time per epoch under different batch sizes and find that training time decreases as the unlearning batch size grows. For instance, on BERT4Rec, the training time per epoch is 0.4 seconds with a batch size of 256, whereas it is 2.5 seconds with a batch size of 32. The model generally converges to satisfactory unlearning performance after approximately 100 epochs. So this demonstrates a trade-off between unlearning effectiveness and efficiency when selecting the unlearning batch size.

### v2

@@ 4.1. Multi-objective unlearning
While effective unlearning is important, maintaining recommendation performance is equally crucial. This necessitates a careful balance between the unlearning objective and the recommendation utility, so we formulate approximate unlearning as a multi-task optimization problem. The key idea is to jointly optimize the recommendation utility and unlearning effectiveness through training a composite objective function with three task-specific losses: (i) an unlearning loss $\mathcal{L}_{unlearn}$ that encourages forgetting specific samples, (ii) a retention loss $\mathcal{L}_{normal}$ to preserve recommendation performance, and (iii) a consistency loss $\mathcal{L}_{KL}$ that regularizes deviations from the previous model.
Next, we introduce the three types of loss function in detail and describe how they are weighted during optimization.
Unlearning loss. Unlearning can be regarded as the inverse process of forward learning, so we apply gradient ascent to explicitly remove the influence of the unlearning sample, which intuitively corresponds to adding back the gradient of the target data that was previously subtracted. For the target item $v_{t}$ to be unlearned, the unlearning loss is formulated as:
(5) $\mathcal{L}_{unlearn}=\log(P_{\theta}(v_{t}|v_{1:t-1})).$
@@
Normal loss. To ensure that the unlearned model retains its general predictive capability, we include a standard cross-entropy loss based on the next item $v_{t+1}$, given the prefix session $\{v_{1},\ldots,v_{t-1}\}$:
(6) $\mathcal{L}_{normal}=-\log(P_{\theta}(v_{t+1}|v_{1:t-1})).$
Since unlearning is performed to remove an item from a session, it is essential to ensure that the model maintains continuity on the remaining items in the session. $\mathcal{L}_{normal}$ guides the model to remain functional in its primary task, i.e., predicting plausible future items based on user behavior, despite forgetting specific items.
@@
(7) $\mathcal{L}_{KL}=KL(P_{\theta^{rec}}(\cdot|v_{1:t-1})||P_{\theta}(\cdot|v_{1:t-1})).$
@@
(8) $\mathcal{L}_{app}=\alpha_{1}\mathcal{L}_{unlearn}+\alpha_{2}\mathcal{L}_{normal}+\alpha_{3}\mathcal{L}_{KL}.$
@@ 4.2. Pareto-optimality solver
A crucial challenge in multi-task optimization is balancing the importance of each objective. Manually tuning the weights $\alpha_{1}$, $\alpha_{2}$, and $\alpha_{3}$ is suboptimal and sensitive to data distribution and model scale. Furthermore, maintaining fixed hyperparameters throughout the entire training process can hinder the model’s ability to adapt to changes and negatively impact performance. To effectively solve this problem, we employ a Pareto-optimality solver, which can reliably find a stable solution that balances the trade-offs between conflicting objectives. This approach ensures that the model simultaneously achieves effective unlearning while maintaining overall recommendation quality. Given a model parameterized by $\theta$ and a series of learning objective functions $\mathcal{L}_{unlearn},\mathcal{L}_{normal},\mathcal{L}_{KL}$, Pareto-optimality is a state in which it is impossible to enhance one objective without compromising another. Formally, Pareto-optimality can be defined as:
@@
Formally, MGDA uses the Karush-Kuhn-Tucker (KKT) conditions (Bertsekas, 1997) to describe constraints on $\alpha_{1},\alpha_{2},\alpha_{3}$:
(10) $\min_{\alpha_{1},\alpha_{2},\alpha_{3}}\left\|\sum_{i=1}^{3}\alpha_{i}\mathbf{g}_{i}\right\|^{2}_{2}\quad\text{subject to}\quad\sum_{i=1}^{3}\alpha_{i}=1,\ \alpha_{i}\geq 0.$
@@
After computing $\alpha_{i}$ for all tasks, we use the loss weighted by $\alpha_{i}$ to update the model parameters, thus identifying an optimization direction that preserves the performance of recommendation while improving the effectiveness of unlearning.
@@ 4.3. Curriculum unlearning
To further enhance effectiveness of approximate unlearning when handling multiple requests simultaneously, we propose a curriculum-based unlearning strategy that prioritizes forgetting easier samples first, progressively moving toward harder ones. The underlying assumption is that removing low-impact easier samples earlier helps stabilize training and facilitates optimization. The curriculum unlearning framework comprises two key components: (i) difficulty estimation and (ii) sample scheduling.
@@ 4.3.1. Difficulty estimation
(11) $\mbox{}\hskip-5.69054pt\mathit{Dif}_{g}(x)\!=\!-\cos\left(\nabla_{\theta}\mathcal{L}_{unlearn}(x),\ \nabla_{\theta}\mathcal{L}_{normal}(x)+\nabla_{\theta}\mathcal{L}_{KL}(x)\right)\!.\hskip-2.84526pt\mbox{}$
@@
Embedding unlearning difficulty. This metric is based on the intuition that samples which the model strongly favors are harder to be forgotten. We compute the dot product between the session state representation and the embedding of the item to be forgotten. Given a session state representation $\mathbf{e}_{session}$ and the embedding of the target item $x$ to be forgotten $\mathbf{e}_{x}$, CAU computes:
(12) $\mathit{Dif}_{e}(x)=\langle\mathbf{e}_{session},\ \mathbf{e}_{x}\rangle.$
Here, $\mathbf{e}_{session}$ is the encoded state representation of the session prefix (e.g., from an RNN or self-attention network), and $\mathbf{e}_{x}$ is the embedding of the target item $x$ to be forgotten. $\langle\rangle$ denotes dot product. A higher $\mathit{Dif}_{e}$ implies the model is more confident in predicting the unlearning target $x$ , making it harder to be removed. Hence, we treat samples with lower $\mathit{Dif}_{e}$ as easier ones and schedule them earlier in the unlearning process.
@@
4   calculate $\mathit{Dif}_{g}$ or $\mathit{Dif}_{e}$ according to Eq. (11) or Eq. (12);
5 $D_{f\text{-sorted}}\leftarrow\text{sort }D_{f}\text{ ascendingly }\text{by $\mathit{Dif}_{g}$ or $\mathit{Dif}_{e}$}$;
@@ 4.3.2. Sample scheduling
For example, when employing the gradient unlearning difficulty metric, the initial phase of each epoch concentrates on samples with high gradient alignment (i.e., lower $\mathit{Dif}_{g}$), indicating that their gradients align well and thus are easier to be optimized. As training progresses, more challenging samples with higher $\mathit{Dif}_{g}$ and more conflicting gradients are incrementally introduced into the training batches. Such curriculum progression reduces abrupt gradient conflicts and helps maintain a stable optimization trajectory.
@@
(13) $p_{x}(t)=\frac{(\exp((2t-1)\cdot(\mathit{Dif}(x)-\overline{\mathit{Dif}})))^{\tau}}{\sum_{j}(\exp((2t-1)\cdot(\mathit{Dif}(j)-\overline{\mathit{Dif}})))^{\tau}},$
where $\mathit{Dif}$ denotes the difficulty scores (i.e., either $\mathit{Dif}_{g}$ or $\mathit{Dif}_{e}$), and $\overline{\mathit{Dif}}$ is the mean difficulty score of all unlearning samples. The time variable $t$ is normalized to the interval $[0,1]$ based on the current training progress, calculated as $t=\mathit{current\_step}/\mathit{total\_step}$. $\tau$ is the temperature coefficient.
@@ 5 $t\leftarrow step/T$;
7      calculate $\mathit{Dif}$ according to Eq. (11) or Eq. (12);
@@ 5.1.1. Datasets and pre-processing
We conduct experiments on three public accessible datasets: MovieLens-1M, [^1] https://grouplens.org/datasets/movielens/1m/ Amazon Beauty and Games. [^2] https://jmcauley.ucsd.edu/data/amazon/ These datasets vary in scale, domain, and sparsity.
@@ 5.1.2. Evaluation protocols
Recommendation performance and unlearning effectiveness follow distinct numerical patterns, such as higher $NDCG@k$ and $Recall@k$ indicate better recommendation performance, while lower $Hit_{u}@k$ signifies better unlearning effectiveness. We therefore introduce a new metric $U_{\beta}$ to facilitate a comprehensive evaluation of both aspects. $U_{\beta}$ integrates $Recall@k$ and $Hit_{u}@k$ as:
@@ 5.1.3. Baselines
  • GRU4Rec (Hidasi et al., 2015) uses gated recurrent units (GRU) to model user interaction sequences which is an RNN-based sequential recommendation model.
@@
  • Retrain: retrain the whole model from scratch on the remaining dataset; this is computationally expensive.
@@ 5.1.4. Implementation details
All methods are learned with the Adam optimizer (Kingma and Ba, 2014). The learning rate is set as 1e31e-3. The model input consists of the most recent 10 interacted items for Beauty, 20 for Games, and 50 for MovieLens-1M, with shorter sequences padded using a special padding token. The training batch size is set to 256, while the unlearning batch size is set to 128. The temperature coefficient $\tau$ is 2. The embedding size is fixed at 64. For the unlearning process in hard-sampling, the unlearning epoch is set to 200 for SASRec and BERT4Rec, and 100 for GRU4Rec. And for soft-sampling, the number of unlearning step is adjusted correspondingly. Each experiment is repeated 5 times, and the average performance is reported.
Table 2. Unlearning effectiveness and overall performance comparison of different methods on the three datasets. The U-score is computed by Recall@10 and $Hit_{u}$@1, with $\beta=10$ for Beauty and $\beta=3$ for Games and ML-1M. Boldface denotes the best score.
@@ 5.2. Recommendation performance
We compare the recommendation performance results in Table 1. We can observe that due to the removal of 10% of the training data, the recommendation performance declines across all baselines, which is an expected outcome. However, in the context of the unlearning task, our primary concern is the ability to preserve recommendation performance on the remaining samples rather than the unlearning effectiveness alone.
From the table, We observe that the recommendation performance of SISA and SRU [^3] SRU results appear lower as we strictly evaluate next-item prediction on the final interaction of each session, whereas the original SRU uses a less restrictive global-sequence average across all sub-sequences. drops markedly, possibly due to the partitioning of the dataset into disjoint shards, which substantially reduces correlations between sessions. Even with the incorporation of attention-based aggregation mechanisms in SRU, its effectiveness in mitigating performance degradation remains limited. Compared with Retrain, their performance drops by 63% and 32%, respectively. Such a drastic reduction is undesirable, as overly low recommendation performance would directly impair the user experience.
In contrast, CAU demonstrates significantly better performance preservation, achieving results close to the Original model with a stable performance drop controlled within 8%. This advantage can be attributed to the introduction of the normal loss and KL loss during the unlearning model update, which suppresses excessive disturbance to the original representation distribution, thereby enabling effective unlearning of target samples while retaining the model’s recommendation capability as much as possible.
@@ 5.3. Unlearning effectiveness and efficiency
Importantly, we can observe that in most cases, soft-sampling is more effective than hard-sampling. For instance, on the ML-1M and BERT4Rec datasets using gradient unlearning difficulty, the U-score of soft-sampling is 0.6913 compared to 0.6711 for hard-sampling. Similarly, with embedding unlearning difficulty, soft-sampling consistently achieves higher scores than hard-sampling. This improvement is likely due to the increased robustness that soft-sampling provides throughout the unlearning procedure.
Table 3. Comparison of unlearning efficiency (minutes, m). Boldface denotes the best score.
@@ 5.4. Ablation study
In this section, we conduct an ablation study to analyze the functionality of the main components and strategies of CAU (i.e., Pareto-optimality solver (PS), curriculum unlearning (CU), and recommendation loss (RL), where RL consists of both the normal loss and the KL loss). Table 4 shows the performance of CAU and its variants on all three datasets using SASRec as the backbone model. Next, we introduce the variants and analyze the effects.
(1) Effect of the Pareto-optimality solver: We removed the Pareto-optimality solver component and assigned an equal weight to each task, which means setting all $\alpha$ values to 0.33. The experimental results demonstrate the effectiveness of the Pareto-optimality solver: by formulating the entire unlearning process as a multi-task optimization problem, the Pareto-optimality solver iteratively searches for gradient update directions that do not degrade any individual sub-task, thereby achieving a better balance between recommendation performance and unlearning effectiveness.
@@
(3) Effect of recommendation loss: We remove both the normal loss and the KL loss, which means directly apply GA to unlearn data. We see that this will lead to catastrophic effects, as the model’s recommendation accuracy almost completely collapses within fewer than five unlearning epochs, and the $Hit_{u}$ value becomes meaningless. This observation highlights the critical role of the recommendation loss in preserving the model’s recommendation capability.
@@ 5.5.1. Impact of unlearning batch size
Next, we investigate the impact of batch size on the unlearning performance during the forgetting process. Because the Pareto-Optimality solver estimates task conflicts and updates the model on a batch basis, the batch size plays a critical role in determining its effectiveness.
Here we choose batch size $\in\{32,64,128,256,512\}$ to conduct experiments. Figure 2 illustrates the performance of the three models on the Games dataset. We clearly observe that as the batch size increases, the U-score tends to decrease. This is because the Pareto-optimality solver dynamically searches for the optimal gradient update on a per-batch basis. And a smaller batch size reduces conflicts among data samples within each batch, increasing the likelihood of identifying accurate update directions, thereby yielding higher U-scores, but may lead to longer training time. We also measure the training time per epoch under different batch sizes and find that training time decreases as the unlearning batch size grows. For instance, on BERT4Rec, the training time per epoch is 0.4 seconds with a batch size of 256, whereas it is 2.5 seconds with a batch size of 32. The model generally converges to satisfactory unlearning performance after approximately 100 epochs. This demonstrates a trade-off between unlearning effectiveness and efficiency when selecting the unlearning batch size.